Determine the number of cubic yards of concrete in a pyramid with a square base measuring 8 yards on each side and an altitude of 42 feet. Express to the nearest hundredth.
Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

Volume = (1/3) Bh
=(1/3)*64*42 ;;;the 64 is the 8*8 area of the base. Volume in cubic yds.
=14*64=896 cubic yds.

Answer by Alan3354(69443)

Determine the number of cubic yards of concrete in a pyramid with a square base measuring 8 yards on each side and an altitude of 42 feet. Express to the nearest hundredth.
--------------
Vol = base area*height/3
= 63*42/3
882 cu yds